Two objects have the same kinetic energy. One has a speed that is 2.1 times greater than the speed of the other. What is the ratio of their masses? (Let m1 be the slower object and m2 be the faster.)

m1/m2 =
Answer #1

here,

the speed of second mass , v2 = 2.1 * speed of first mass

v2 = 2.1 * v1

as they have same kinetic energy

KE1 = KE2

0.5 * m1 * v1^2 = 0.5 * m2 * v2^2

m1 /m2 = (v2 / v1)^2

m1 /m2 = ( 2.1 v1 /v1)^2

m1 /m2 = 4.41

the ratio of masses is 4.41